OpenStudy (anonymous):

a quick question about working with logs is this true? ln(x) + ln(y) = ln(x + y)

OpenStudy (agent0smith):

No. loga+logb = logab

OpenStudy (isaiah.feynman):

\[\ln (a)+\ln (b)=\ln (ab)\]

OpenStudy (anonymous):

so if I have ln(x +y), is there any way to simplify that, or do I just work with it as it is?

OpenStudy (isaiah.feynman):

Yes. For example if you have \[\ln (5+9) \] that equals \[\ln (14)\]. But if you have \[\ln (5)+\ln (9)\] that equals \[\ln (5\times9) = \ln (45)\]

OpenStudy (anonymous):

ok, that makes sense, thank you. I'm actually working on differentation and forgot the log rules. what I actually have is ln(x^2 + x + 1) so I was looking for the rule that would apply. Since I don't know the value of my x I'll leave it as it is or use substitution. Thank you.

OpenStudy (isaiah.feynman):

Wait are you trying to differentiate that function?

OpenStudy (anonymous):

it's using logaritmic differentiaton to find the derivitive. I"m using the online derivative calculator, but making sure I understand all the steps too. I"m sort of taking calc 2 10 years after I took calc 1. Having some difficutlies

OpenStudy (anonymous):

my internet is cutting in and out so I'll keep plugging away at it

OpenStudy (isaiah.feynman):

See. \[f(x)=\ln (x^{2}+x+1)\]\[f \prime(x)=\frac{ 1 }{ x^{2}+x+1 }\times(x^{2}+x+1)\prime \]

OpenStudy (anonymous):

Chain rule!!! I am getting some of it. thank you

OpenStudy (agent0smith):

Derivative of ln(f(x)) is \[\Large \frac{ f'(x) }{ f(x)}\] or the derivative of the function insde the log, divided by the function inside the log